# A Huber Loss Minimization Approach to Mean Estimation under User-level Differential Privacy

### v1

@@ 1 Introduction
Our method has the following advantages. Firstly, our method adapts well to imbalanced datasets, since the threshold $T_{i}$ of Huber loss are selected adaptively according to the sample size per user, which leads to a better tradeoff between sensitivity and bias. Secondly, our method performs better for heavy-tailed distributions, since we control sensitivity by penalizing large distances using Huber loss, which yields a smaller bias than the clipping operation. Apart from solving these practical issues, it worths mentioning that our method solves robustness and privacy issues simultaneously. In modern data analysis, it is common for a system to suffer from both poisoning and inference attacks at the same time [41]. Consequently, many recent works focus on unified methods for item-level DP and robustness to cope with both attacks simultaneously [42, 43, 44, 45]. To the best of our knowledge, our method is the first attempt to unify robustness and DP at user-level.
@@ 2 Related Work
User-level DP. [22] applies a brute-force clipping method [17] for user-level DP in federated learning. [27] made the first step towards optimal rates under user-level DP, which analyzed discrete distribution estimation problems. The popular method WME was proposed in [28], which uses the idea of two-stage approaches [46, 47, 45]. The two-stage method for user-level DP has also been extended to stochastic optimization problems [48, 49]. [50] analyzes mean estimation for boolean signals under user-level DP in heterogeneous settings. There are also some works focusing on black-box conversion from item-level DP to user-level counterparts. [29] analyzed general statistical problems, which shows that a class of algorithms for item-level DP problems having the pseudo-globally stable property can be converted into user-level DP algorithms. Following [29], [51] expanded such transformation for any item-level algorithms. [30] extends the works to smaller $m$. It is discussed in [49] that these black-box methods have suboptimal dependence on $\epsilon$.
From robustness to DP. Robustness and DP have close relationships since they both require the outputs to be insensitive to minor changes in input samples. There are three types of methods for conversion from robust statistics to DP. The first one is propose-test-release (PTR), which was first proposed in [52], and was extended into high dimensional cases in [53]. The second choice is smooth sensitivity [40], which calculates the noise based on the "smoothed" local sensitivity. For example, [54] designed a method to protect trimmed mean with smooth sensitivity. The third solution is inverse sensitivity [55, 56, 5], which can achieve pure differential privacy (i.e. $\delta=0$). All these methods require a detailed analysis of the sensitivity. For some recently proposed high dimensional estimators [57, 58, 59], the sensitivity is usually large and hard to analyze. As a common method for robust statistics [39], Huber loss minimization has been widely applied in robust regression [60, 61], denoising [62] and robust federated learning [63]. To the best of our knowledge, Huber loss minimization has not been applied to DP.
Our work is the first attempt to use the Huber loss minimization method in user-level DP. With an adaptive selection of weights and connecting points between quadratic and linear parts, our method achieves better performance for imbalanced users and heavy-tailed distributions. Existing analysis on the robustness of the Huber loss minimizer can not be directly used here since we need to bound the sensitivity for all possible datasets. The smooth sensitivity framework [40] is applied here to determine the strength of additive noise for privacy protection.
@@ 5 Analysis: Balanced Users
$G(\mathcal{D},k)\leq\frac{2T}{n-k-\Delta(\mathcal{D})};$ (16)
@@ 5.1 Bounded Support
For the simplest case with bounded support and balanced users, the two-stage approach in [28] is already nearly optimal (Corollary 1 in [28]). Therefore, improvement in polynomial factors is impossible. Nevertheless, we still improve on the logarithm factor.
@@ 6 Analysis: Imbalanced Users
Then the convergence of mean squared error is shown in Theorem 5.
@@ 7.2 Imbalanced Users
Now we show the performance with unbalanced users. For some $\gamma\geq 1$, let $s_{i}=\lceil N(i/n)^{\gamma}\rceil$ for $i=0,\ldots,n$, and $m_{i}=s_{i}-s_{i-1}$ for $i=1,\ldots,n$. It can be shown that Assumption 3 is satisfied with $\gamma$. Therefore, according to the analysis in Section 6, we let $w_{i}=m_{i}\wedge m_{c}/\sum_{j}m_{j}\wedge m_{c}$, with $m_{c}=\gamma N/n$, and $T_{i}=A/\sqrt{m_{i}\wedge m_{c}}$, in which $A$ is tuned optimally for each case. Figure 2 shows the growth curve of mean squared error with respect to $\gamma$.
@@ Appendix I Proof of Lemma 4
$\bar{\mathbf{X}^{*}}=\left\{\begin{array}[]{ccc}\bar{\mathbf{X}}&\text{if}&\left\lVert\bar{\mathbf{X}}-\mu\right\rVert\leq r_{0}\\ \mu&\text{if}&\left\lVert\bar{\mathbf{X}}-\mu\right\rVert\leq r_{0},\end{array}\right.$ (173)

### v2

@@ 1 Introduction
Our method has the following advantages. Firstly, our method adapts well to imbalanced datasets, since the threshold $T_{i}$ of Huber loss are selected adaptively according to the sample size per user, which leads to a better tradeoff between sensitivity and bias. Secondly, our method performs better for heavy-tailed distributions, since we control sensitivity by penalizing large distances using Huber loss, which yields a smaller bias than the clipping operation. Apart from solving these practical issues, it worths mentioning that our method solves robustness (to model poisoning attacks) and privacy issues simultaneously. In modern data analysis, it is common for a system to suffer from both poisoning and inference attacks at the same time [41, 42, 43]. Consequently, many recent works focus on unified methods for item-level DP and robustness to cope with both attacks simultaneously [44, 45, 46, 47]. To the best of our knowledge, our method is the first attempt to unify robustness and DP at user-level.
@@ 2 Related Work
User-level DP. [22] applies a brute-force clipping method [17] for user-level DP in federated learning. [27] made the first step towards optimal rates under user-level DP, which analyzed discrete distribution estimation problems. The popular method WME was proposed in [28], which uses the idea of two-stage approaches [48, 49, 47]. The two-stage method for user-level DP has also been extended to stochastic optimization problems [50, 51]. [52] analyzes mean estimation for boolean signals under user-level DP in heterogeneous settings. There are also some works focusing on black-box conversion from item-level DP to user-level counterparts. [29] analyzed general statistical problems, which shows that a class of algorithms for item-level DP problems having the pseudo-globally stable property can be converted into user-level DP algorithms. Following [29], [53] expanded such transformation for any item-level algorithms. [30] extends the works to smaller $m$. It is discussed in [51] that these black-box methods have suboptimal dependence on $\epsilon$. [54, 55, 56] studies user-level DP under local model.
From robustness to DP. Robustness and DP have close relationships since they both require the outputs to be insensitive to minor changes in input samples. There are three types of methods for conversion from robust statistics to DP. The first one is propose-test-release (PTR), which was first proposed in [57], and was extended into high dimensional cases in [58]. The second choice is smooth sensitivity [40], which calculates the noise based on the "smoothed" local sensitivity. For example, [59] designed a method to protect trimmed mean with smooth sensitivity. The third solution is inverse sensitivity [60, 61, 5], which can achieve pure differential privacy (i.e. $\delta=0$). All these methods require a detailed analysis of the sensitivity. For some recently proposed high dimensional estimators [62, 63, 64], the sensitivity is usually large and hard to analyze. As a common method for robust statistics [39], Huber loss minimization has been widely applied in robust regression [65, 66], denoising [67] and robust federated learning [68]. Huber loss has also been used in DP [69, 70] for linear regression problems.
Concurrent work. After the initial submission of this paper, we notice an independent work [71], which also studies mean estimation under user-level DP (which is called person-level DP in [71]). [71] considers directional bound, which requires that the moment is bounded in every direction. However, we consider non-directional bound, which bounds the $\ell_{2}$ norm of a random vector. We refer to Section 1.3.1 in [71] for further discussion.
Our work is the first attempt to use the Huber loss minimization method in user-level DP. With an adaptive selection of weights and connecting points between quadratic and linear parts, our method achieves a significantly better performance for imbalanced users and heavy-tailed distributions.
@@ 5 Analysis: Balanced Users
$G(\mathcal{D},k)=\frac{2T}{n-k-\Delta(\mathcal{D})};$ (16)
@@ 5.1 Bounded Support
For the simplest case with bounded support and balanced users, the two-stage approach in [28] is already nearly optimal (Corollary 1 in [28]). Therefore, improvement in polynomial factors is impossible. Nevertheless, we still improve on the logarithm factor. The main purpose of Theorem 2 is to show that our improvement on heavy-tailed distributions and imbalanced users is not at the cost of hurting the performance under the simplest case with bounded distributions and balanced users.
@@ 6 Analysis: Imbalanced Users
In Assumption 3, $\gamma$ can be viewed as the degree of imbalance. For a better explanation, we provide the following examples:
  • If users are balanced, then $\gamma=1$;
  • If the $i$-th user has $ki$ samples (which means that the number of items belonging to each user is linear in its order), then for large $n$, $\gamma$ is approximately $\sqrt{2}$.
In general, $\gamma$ is large if users are highly imbalanced. Under Assumption 3, the convergence of mean squared error is shown in Theorem 5.
@@ 7.2 Imbalanced Users
Now we show the performance with unbalanced users. For some $\gamma\geq 1$, let $s_{i}=\lceil N(i/n)^{\gamma}\rceil$ for $i=0,\ldots,n$, and $m_{i}=s_{i}-s_{i-1}$ for $i=1,\ldots,n$. It can be shown that Assumption 3 is satisfied with $\gamma$. Therefore, according to the analysis in Section 6, we let $w_{i}=m_{i}\wedge m_{c}/\sum_{j}m_{j}\wedge m_{c}$, with $m_{c}=\gamma N/n$, and $T_{i}=A/\sqrt{m_{i}\wedge m_{c}}$, in which $A$ is tuned optimally for each case. The selection of $T_{i}$ may be slightly different from (25) in Theorem 5. In Theorem 5, $T_{i}$ is selected to minimize the theoretical upper bound. To ensure that the analysis is mathematically rigorous, the upper bound of estimation error is larger than the truth. Therefore, the optimal value of $T_{i}$ in practice is slightly different from that derived in theories. Note that such parameter tuning does not require additional privacy budget since in each experiment, $T_{i}$ are hyperparameters that is fixed before knowing the value of each sample. They are not determined adaptively based on the data. Figure 2 shows the growth curve of mean squared error with respect to $\gamma$.
@@ Appendix I Proof of Lemma 4
$\bar{\mathbf{X}^{*}}=\left\{\begin{array}[]{ccc}\bar{\mathbf{X}}&\text{if}&\left\lVert\bar{\mathbf{X}}-\mu\right\rVert\leq r_{0}\\ \mu&\text{if}&\left\lVert\bar{\mathbf{X}}-\mu\right\rVert>r_{0},\end{array}\right.$ (173)